OpenStudy (anonymous):

how to solve lim x->0 sin4x/x

OpenStudy (hba):

Multiply and divide by 4

OpenStudy (anonymous):

how to solve can you give steps

OpenStudy (hba):

Yes multiply and divide by four first.

OpenStudy (hba):

@Ghayyas Do it.

OpenStudy (anonymous):

can you give me full steps

OpenStudy (hba):

@Ghayyas That is all what you have to do.

OpenStudy (hba):

Can you do it and show that to me.

OpenStudy (anonymous):

hba how?

OpenStudy (hba):

\[\lim_{x \rightarrow 0} \frac{ \sin4x }{ x } \times \frac{ 4 }{ 4 }\]

OpenStudy (hba):

Like this.

OpenStudy (hba):

Do you know that Sin 4x/4x=1 If you need a proof please check my proof here. http://openstudy.com/users/hba#/updates/50e3edefe4b028291d74cd6f

OpenStudy (hba):

Now apply limit on this \[\lim_{x \rightarrow 0} \frac{ \sin4x }{4 x } \times \ 4\]

OpenStudy (hba):

What will you get ?

OpenStudy (anonymous):

but in my book answer is 4

OpenStudy (hba):

Yes apply limit

OpenStudy (anonymous):

how to solve this?